OpenStudy (anonymous):

solve Solve. 2^(x + 3) = 7

OpenStudy (anonymous):

\[\LARGE 2^{x+3}=7 \Longrightarrow \log(2)^{x +3}=\log(7)\] \[\LARGE (x+3)\cdot \log2=\log7\] \[\LARGE x+3=\frac{\log7}{\log2}\] add -3 to each side... and go grab a calculator !!

OpenStudy (anonymous):

THANK YOU TUPAC

OpenStudy (anonymous):

@kres but look tupac it makes no sense

OpenStudy (experimentx):

use calc for the rest best one is here http://www.wolframalpha.com

OpenStudy (anonymous):

i did but it dsnt show me what i need

OpenStudy (anonymous):

oh i put it in backwards! thank you

OpenStudy (anonymous):

so it was -.19

OpenStudy (experimentx):

yeah //

OpenStudy (anonymous):

@cheeto365 is that a test or something? .. Online? where is it?

OpenStudy (anonymous):

its a pretest for my math class online its just confusing cause of the unequal bases